Problem: An administrator of a popular web site is told that a new server can handle 41,000 "hits" per second. The web site currently experiences a peak demand of about 105,000 hits per second; but every month, the peak demand increases by 2800 hits per second. Use a proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 24 months.
Answer by stanbon(48568)

Problem: An administrator of a popular web site is told that a new server can handle 41,000 "hits" per second. The web site currently experiences a peak demand of about 105,000 hits per second; but every month, the peak demand increases by 2800 hits per second. Use a proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 24 months.
-----------------------
The function is linear, not proportional.
---------------------
slope = 2800 hit/month
Intercept = 105,000
-------------------
N(x) = 2800x + 105000
N(24) = 2800*24 + 105000
N(24) = 67200 + 105000
N(24) = 172200 bits/second
172200/41000 = 4.2 ; rounding up gives 5
(5 is the # of servers he will need in 24 months)
=================
Cheers,
Stan H.